- Press Release
- Dec 8, 2022
NASA Space Station On-Orbit Status 3 Mar 2004
All ISS systems continue to function nominally, except as noted previously or below.
Before breakfast both crewmembers completed another session of the periodic Russian medical experiment protocols PZEh-MO-7 (calf volume measurement) and PZEh-MO-8 (body mass measurement). FE Alexander Kaleri set up the MO-8 “scales” equipment and later broke it down and stowed it away. (Last time done: 2/16). [For operational monitoring and data recording the special IFEP (in-flight examination program) software is used on the medical equipment computer (MEC).]
On the MSG (Microgravity Science Glovebox), CDR Michael Foale terminated sample processing of the PromISS-3 experiment (Protein Crystal Growth Monitoring by Digital Holographic Microscope #3). Later, after a 1-hr. delay to allow the sample wheel to return to the zero position, Foale removed the payload from the MSG. The glovebox rack was then powered down.
Alex Kaleri activated the Aquarius-B (AQUB-01) payload hardware in the Service Module (SM) and transferred the PromISS-3 samples from the MSG inside their return container in a Ziploc bag to the AQUB-01 incubator.
The FE performed another daily session of inspecting the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse. [Maintenance of the experiment involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording.]
Sasha also activated the SKR-05 hardware of the Russian TEKh-25 Skorpio (“scorpion”) payload in the SM “floor” underneath the TVIS, which he had powered down on 2/24. [Skorpio’s objective is to monitor environmental radiation parameters with dosimeters inside station compartments at various places and to characterize environmental conditions for conducting scientific and technical experiments. To operate, Skorpio requires about 6 W of energy.]
The crew performed periodic station air sampling. Alex first tested for CO (carbon monoxide) levels in SM and DC-1 by collecting samples with the IPD Draeger tubes; then he tested the SM for Freon (coolant used in the SKV air conditioner) with the AK-1M-F. Mike meanwhile “grabbed” air samples with the GSC (grab sample container) at the center of the SM, then used the new Dual Sorbent Tube (DST), instead of the SSAS (Solid Sorbent Air Sampler), to take samples in the center of the Lab and SM.
Kaleri began a check of communication between the Russian Wiener power laptop and the BSPN server of the Matryoshka payload equipment. Today, as Part 1, he performed time synchronization between the two units, supported by tagup with a ground specialist. [Tomorrow’s Part 2 will be to copy software files from the server to a flash card via the laptop.]
Foale continued transferring those equipment items back to the U.S. segment (USOS) that were brought to the Russian segment (RS) for the duration of the EVA-9 isolation.
Sasha returned the items, which he had stocked in the crew’s EVA medical kit on 2/24, back to their original containers and locations.
On the HRF (Human Research Facility) rack, the CDR reconfigured laptop settings and installed a number of audio drivers to support upcoming ADUM OPE (advanced ultrasound on-board proficiency enhancer) operations. The PC, connected to the HRF via Ethernet, was left on to allow the ground repeating downlinking a file with stored data of the FOOT (Foot/Ground Reaction Forces during Space Flight) experiment. [An earlier attempt at downlink had failed. The new downlink required two hours, and the PC was to be powered down after ground receipt was confirmed.]
The FE terminated the discharge of the first Orlan spacesuit battery (825M3) in the ZU-S charging unit, then installed the second battery for its discharge.
Sasha also prepared the daily IMS (inventory management system) delta file for automatic export/import to update the databases.
Last night’s station reboost took place as planned (5:59pm EST) and was successful. The one-burn maneuver of ~2.3 m/s delta-V raised apogee by 7.6 km, boosting mean orbit altitude by 3.8 km. Thus, eccentricity (“ellipticity”) of the orbit also went up. After the reboost, station attitude was maneuvered to solar-oriented XPOP (x-axis perpendicular to orbit plane)
Lab window constraints while in XPOP attitudes still are in effect. [They require the window shutter to be open for a cumulative period no longer than 1 hr in a 24-hr period. Also, while in XPOP, for thermal reasons the window shutter can only be opened for a maximum of 15 min at a time; prior to re-opening it must be kept closed for the amount of time it was open plus 5 min.While in nominal LVLH (local vertical/local horizontal) XVV or YVV attitudes, the window shuttermay be open for a cumulative periodno longer than 1.5 hours in a 24-hour period. After a successful window evacuation on Friday (3/5)and at least two pressure checks of the inter-pane “Volume B”, these window constraints will be re-evaluated.]
Today’s CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Lake Nasser, Toshka Lakes, Egypt (nadir views of the new lakes at Toshka: shooting detailed images of the margins of the new lake nearest the Nile where there is much activity in road building, canal building, and laying out of agricultural plots), Lake Chad, Chad (long-term monitoring site. Until the late 60s, Lake Chad water levels were consistently high. They have been consistently low since, with a slight trend to filling. Crew was to shoot a status image. Chad levels are interesting since they indicate some multi-El Nino cyclicity that is not understood), and Bamako, Mali (nadir pass over this city situated in the Niger inland delta).
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at
U.S. and Russian Segment Status (as of today, 2:30pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is Off. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 26.4; ppO2 (mmHg) — 150.9; ppCO2 (mmHg) — 3.8;
- SM Transfer Compartment: Pressure (mmHg) — 741; temperature (deg C) — 19.8.
- FGB Cabin: Pressure (mmHg) — 744; temperature (deg C) — 22.7.
- Node: Pressure (mmHg) — 745.80; temperature (deg C) — 23.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 747.24; temperature (deg C) — 24.9; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 747.44; temperature (deg C) — 23.3; shell heater temp (deg C) — 23.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- PMA-1: Shell heater temp (deg C) — 25.1
- PMA-2: Shell heater temp (deg C) — 13.9
(n/a = data not available)
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in in Autotrack (suntracking) and bias-angled 43 deg. for drag reduction.
- SM batteries: All batteries (8) are in “Partial Charge” mode.
- FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup.
- INT-2 is operating; INT-1 is Off.
- EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).
- Total propellant load available: 4036 kg (8898 lb) as of 2/27/04 [SM(755) + FGB(2622) + Progress M(0) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 3 CMGs on-line (CMG-1 failed).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
- XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist).
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- *S-band is operating nominally (on string 2).
- Ku-band is operating nominally.
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, powered on both strings.
- MBS: KA power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 5:16am EST [= epoch]):
- Mean altitude — 369.1 km
- Apogee — 376.8km
- Perigee — 361.5 km
- Period — 91.93 min.
- Inclination (to Equator) — 51.628 deg
- Eccentricity — 0.0011325
- Orbits per 24-hr. day — 15.67
- Mean altitude gain last 24 hours — 3800 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 30172
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html